OpenStudy (anonymous):

Find the normal form of the equation of the plane that passes through P = (0,-2,5) and is parallel to the plane with general equation 6x-y+2z = 3

OpenStudy (turingtest):

if two planes are parallel they have the same normal vectors agree?

OpenStudy (anonymous):

Yes.

OpenStudy (turingtest):

and what is that vector?

OpenStudy (anonymous):

The normal vector would be perpendicular to the line in the question.

OpenStudy (turingtest):

what line in the question? there are only two planes...

OpenStudy (anonymous):

The normal vector is perpendicular to the planes.

OpenStudy (turingtest):

right, and what is it?

OpenStudy (anonymous):

n (dot) x = n (dot) p

OpenStudy (turingtest):

I mean you need to be able to explicitly tell me what n is or you cannot proceed

OpenStudy (turingtest):

for an equation of the form\[ax+by+cz=d\]the normal vector is\[\vec n=\langle a,b,c\rangle\]

OpenStudy (anonymous):

6x - y +2z = 12

OpenStudy (anonymous):

Is that right?

OpenStudy (turingtest):

the normal vector should just be the coefficients of the variables, so no

OpenStudy (anonymous):

The normal vector would just be 6 - 1 + 2.

OpenStudy (turingtest):

n=(6,-1,2) yes

OpenStudy (turingtest):

now you can use your formula n (dot) x = n (dot) p

OpenStudy (turingtest):

I think you have the same lame teacher as someone else I just helped with a similar problem who was using similar notation :P

OpenStudy (anonymous):

Well that helps.

OpenStudy (turingtest):

do you know how to use that formula?

OpenStudy (anonymous):

6 x 6 0 -1(dot) y = -1 (dot) -2 2 z 2 5

OpenStudy (turingtest):

good, which is what?

OpenStudy (anonymous):

<6x, -y, 2z> = 12

OpenStudy (turingtest):

that should not be a vector after taking the dot product; the dot product produces scalars only

OpenStudy (turingtest):

the components need to be added together

OpenStudy (turingtest):

...talking about the left side...

OpenStudy (anonymous):

6x-1y+2z=12

OpenStudy (turingtest):

very good, and you're done!

OpenStudy (anonymous):

Thank you!

OpenStudy (turingtest):

welcome!